OpenStudy (anonymous):

What is the vertex of this quadratic function? f(x) = 2x^2 + 4x – 5

OpenStudy (harsimran_hs4):

want to solve it graphically or algebrically?

Parth (parthkohli):

\(\textbf{HINT}\) Complete the square.

OpenStudy (anonymous):

would it be (-1,-7)?

OpenStudy (anonymous):

The vertex of the quadratic function \(\large f(x)=ax^2+bx+c \) occur at (h, k) where h=-b/(2a) k=f(h)

Parth (parthkohli):

Indeed, that's another way.

OpenStudy (anonymous):

complete the square to get your equation into vertex form: \(y = a(x – h)^2 + k\) or you could do \(x=-{b\over2a}\) to find the x value of the vertex and then plug in the \(x\) into the function to get he y-value

OpenStudy (anonymous):

@venom936 , yes your answer is correct